OpenStudy (anonymous):

two bodies of mass ma and mb are dropped from two different hieghts a and b.the ratio times taken by the two bodies to drop through these distances is

OpenStudy (anonymous):

Tell me if it was helpful or not. thank u.

OpenStudy (anonymous):

for a body of mass m dropped from a height h the time taken is independent of its mass it is given by \[t \alpha c \sqrt{h/g}\] since g is gravitational constant and c is a constnt of proportionnality hence\[t \alpha \sqrt{h}\] hence your answer will be \[t _{1}/t _{2}=\sqrt{a/b}\]

OpenStudy (anonymous):

the ratio is 1/r squared....r is the distance